### v2

# On the well-posedness of SPDEs with locally Lipschitz coefficients

Abstract.
We consider the stochastic partial differential equation,
$\partial_{t}u=\tfrac{1}{2}\partial^{2}_{x}u+b(u)+\sigma(u)\dot{W},$
where $u=u(t\,,x)$ is defined for $(t\,,x)\in(0\,,\infty)\times\mathbb{R}$, and $\dot{W}$ denotes space-time white noise. We prove that this SPDE is well posed solely under the assumptions that the initial condition $u(0)$ is bounded and measurable, and $b$ and $\sigma$ are locally Lipschitz continuous functions and have at most linear growth. Our method is based on a truncation argument together with moment bounds and tail estimates of the truncated solution. The results naturally generalize to the case where $b$ and $\sigma$ are time dependent with uniform-in-time growth and oscillation properties. Additionally, our method can be extended to the stochastic wave equation.
Keywords: SPDEs, space-time white noise, existence and uniqueness.
AMS 2010 subject classification: 60H15; 60H07, 60F05.
The research of the second author is supported in part by the United States’ National Science Foundation grants DMS-1855439 and DMS-2245242
The third author acknowledges support from the Spanish MINECO grant PID2022-138268NB-100 and Ayudas Fundacion BBVA a Proyectos de Investigación Científica 2021
1. Introduction
We revisit the well-posedness of the solution $u=\{u(t\,,x)\}_{t\geq 0,x\in\mathbb{R}}$ of the following stochastic PDE (SPDE):
$\partial_{t}u(t\,,x)=\tfrac{1}{2}\partial^{2}_{x}u(t\,,x)+b(t\,,u(t\,,x))+\sigma(t\,,u(t\,,x))\dot{W}(t\,,x),$ (1.1)
where $(t\,,x)\in(0\,,\infty)\times\mathbb{R}$, subject to $u(0\,,x)=u_{0}(x)$, and the forcing $\dot{W}$ is space-time white noise; that is, $\dot{W}$ is a generalized Gaussian random field with mean zero and $\operatorname{\mathrm{Cov}}[\dot{W}(t\,,x)\,,\dot{W}(s\,,y)]=\delta_{0}(t-s)\delta_{0}(x-y)$ for all $t,s\geq 0$ and $x,y\in\mathbb{R}$. It is well known that (1.1) is well posed when $b$ and $\sigma$ are Lipschitz in their spatial variable uniformly in their time variable; see for example Dalang [Dalang] and Walsh [Walsh]. We aim to extend this result to the setting where “Lipschitz” is replaced by “local Lipschitz with linear growth.”
The present undertaking yields an infinite-dimensional version of one of the foundational results of the theory of stochastic differential equations (SDEs) which asserts that if $Y=\{Y_{t}\}_{t\geq 0}$ denotes a standard Brownian motion on $\mathbb{R}$ and $x_{0}\in\mathbb{R}$ is non random, then the one-dimensional Itô-type SDE $\mathrm{d}X_{t}=b(t\,,X_{t})\,\mathrm{d}t+\sigma(t\,,X_{t})\,\mathrm{d}Y_{t}$ ($t>0$) subject to $X_{0}=x_{0}$ has a unique solution provided only that $b(t\,,x)$ and $\sigma(t\,,x)$ are locally Lipschitz in $x$ with at-most linear growth, all valid uniformly in $t$. The preceding follows from the more classical result about SDEs with Lipschitz coefficients and a stopping time argument. See for example Exercise (2.10) of Revuz and Yor [RevuzYor, p. 383]. It is possible to use essentially the same method in order to extend the preceding to the multidimensional setting where $Y$ denotes a $d$-dimensional Brownian motion, $b:(0\,,\infty)\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\sigma:(0\,,\infty)\times\mathbb{R}^{d}\to(\mathbb{R}^{d})^{2}$.
An infinite-dimensional extension appears in a forthcoming monograph by Robert C. Dalang and Marta Sanz-Solé [bookDS]. There, Dalang and Sanz-Solé show the well-posedness of (1.1), where instead $(t\,,x)\in(0\,,\infty)\times I$ in the case that $I\subset\mathbb{R}$ is a bounded interval (together with any of the usual boundary conditions). A key step in their analysis is the observation that if $b$ and $\sigma$ are locally Lipschitz in their space variable, uniformly in the time variable, then (1.3) below has a predictable random-field solution up to the stopping time $\tau(I)=\inf\{t>0:\sup_{x\in I}|u(t\,,x)|=\infty\}$ $[\inf\varnothing=\infty]$, and $\mathrm{P}\{\tau(I)>0\}=1$ when $I\subset\mathbb{R}$ is bounded. Such stopping-time arguments are typically not used to study equations on $\mathbb{R}_{+}\times\mathbb{R}$ because one expects the resulting solution, if there is one, to be unbounded with probability one. Therefore, in order to produce solutions to the SPDE (1.1) one needs to introduce a different argument.
The purpose of this article is to provide one such argument: We use truncation, as has been done previously, but replace the stopping-time argument by pointwise tail probability estimates for the truncated solution. In this way we are able to show that, with high probability, the truncated solution is not large, uniformly in the truncation level. At that stage, our method boils down to an application of an amusing elementary lemma on summability [Lemma 7] that plays an essential role in this work.
It is also worth noting that the techniques developed in this paper extend beyond the stochastic heat equation (1.1). Specifically, similar methods can be applied to the stochastic wave equation of the following form,
$\partial^{2}_{t}u(t\,,x)=\tfrac{1}{2}\partial^{2}_{x}u(t\,,x)+b(t\,,u(t\,,x))+\sigma(t\,,u(t\,,x))\dot{W}(t\,,x),$
where $t\geq 0$, $x\in\mathbb{R}$, subject to appropriate initial conditions, see [bookDS, Chapter 4] for the existence and uniqueness under globally Lipschitz condition. Although the stochastic wave equation presents additional challenges due to its hyperbolic nature, the local Lipschitz and linear growth assumptions on $b$ and $\sigma$ allow our approach to handle the existence and uniqueness of solutions in this setting as well. The key adjustments consist of utilizing the wave kernel in place of the heat kernel and establishing similar tail probability estimates for the truncated solutions.
We impose the following assumptions on the initial profile $u_{0}$ and the coefficients $b$ and $\sigma$ in (1.1):
Assumption 1.
$u_{0}:\mathbb{R}\to\mathbb{R}$ is non-random, bounded, and measurable.
Assumption 2.
The functions $b:(0\,,\infty)\times\mathbb{R}\rightarrow\mathbb{R}$ and $\sigma:(0\,,\infty)\times\mathbb{R}\rightarrow\mathbb{R}$ are locally Lipschitz continuous in their space variable with at most linear growth, uniformly in their time variable. In other words, $0<\text{\rm Lip}_{n}(b),\text{\rm Lip}_{n}(\sigma)<\infty$ and $0\leq L_{b},L_{\sigma}<\infty,$ for all real numbers $n>0$ where, for every space-time function $\psi$,
$\begin{split}L_{\psi}&=\sup_{t>0}\sup_{x\in\mathbb{R}}\frac{|\psi(t\,,x)|}{1+|x|},\ \text{\rm Lip}_{n}(\psi)=\sup_{t>0}\sup_{\begin{subarray}{c}x,y\in[-n,n]\\ x\neq y\end{subarray}}\frac{|\psi(t\,,x)-\psi(t\,,y)|}{|y-x|}.\end{split}$ (1.2)
Let us recall that a random field solution to (1.1) is a predictable random field $u=\{u(t\,,x)\}_{t\geq 0,x\in\mathbb{R}}$ that satisfies the following integral equation:
$u(t\,,x)=(p_{t}*u_{0})(x)+\mathcal{I}_{b}(t\,,x)+\mathcal{I}_{\sigma}(t\,,x),$ (1.3)
where, for every space-time function $\psi$,
$\mathcal{I}_{\psi}(t\,,x)=\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\psi(s\,,u(s\,,y))\,\mathrm{d}s\,\mathrm{d}y,$ (1.4)
the symbol “*” denotes convolution, and
$p_{r}(z)=\frac{\exp\{-z^{2}/(2r)\}}{\sqrt{2\pi r}}\qquad\text{for all $r>0$ and $z\in\mathbb{R}$}.$
The following is the main result of this paper.
Theorem 3.
Under Assumptions 1 and 2, (1.1) has a unique random-field solution that satisfies the following:
$\adjustlimits{\sup}_{t\in[0,T]}{\sup}_{x\in\mathbb{R}}\mathrm{E}\left(|u(t\,,x)|^{k}\right)<\infty\quad\text{for all $T>0$ and $k\geq 1$}.$
Let us end the Introduction with a brief description outline of the paper. In §2 we introduce the truncated solution and develop some moment and tail estimates. The remaining details of the proof of Theorem 3 are gathered in §3, and use the earlier results of the paper.
Throughout this paper, we write $\|X\|_{p}=\{\mathrm{E}(|X|^{p})\}^{1/p}$ for all $p\geq 1$ and $X\in L^{p}(\Omega)$. For every space-time function $f:(0\,,\infty)\times\mathbb{R}\to\mathbb{R}$, $\text{\rm Lip}(f)$ denotes the optimal Lipschitz constant of $f$; that is,
$\text{\rm Lip}(f)={\color[rgb]{0,0,0}\sup_{t>0}\sup_{\begin{subarray}{c}a,b\in\mathbb{R}\\ a\neq b\end{subarray}}\frac{|f(t\,,b)-f(t\,,a)|}{|b-a|}.}$ (1.5)
In particular, $f$ is globally Lipschitz continuous in $x$, uniformly in $t$, iff $\text{\rm Lip}(f)<\infty$. If $f$ depends only on a spatial variable $x$, then $\text{\rm Lip}(f)$ still makes sense provided that we extend $f$ to a space-time function as follows $f(t\,,x)=f(x)$, in the usual way.
2. Truncation, and preliminary results
In this section we describe the truncation argument that was referred to verbally in the Introduction. Define,
$L_{N,b}=\text{\rm Lip}_{\exp(N)}(b)\quad\text{ and }\quad L_{N,\sigma}=\text{\rm Lip}_{\exp(N)}(\sigma),$ (2.1)
and, for every real number $N>0$, two functions $b_{N}:(0\,,\infty)\times\mathbb{R}\to\mathbb{R}$ and $\sigma_{N}:(0\,,\infty)\times\mathbb{R}\to\mathbb{R}$ as follows: For all $t>0$ and $\psi:\mathbb{R}_{+}\times\mathbb{R}\to\mathbb{R}$ let
$\psi_{N}(t\,,x)=\psi(t\,,x){\bf 1}_{\{-\mathrm{e}^{N}\leq x\leq\mathrm{e}^{N}\}}+\psi(t\,,\mathrm{e}^{N}){\bf 1}_{\{x>\mathrm{e}^{N}\}}++\psi(t\,,-\mathrm{e}^{N}){\bf 1}_{\{x<-\mathrm{e}^{N}\}}.$
Choose and fix an arbitrary $N>0$ and note that $b_{N}$ and $\sigma_{N}$ are globally Lipschitz. In fact, recall (1.5) in order to see that $\text{\rm Lip}(b_{N})=L_{N,b}$ and $\text{\rm Lip}(\sigma_{N})=L_{N,\sigma}$, and
$\adjustlimits{\sup}_{t>0}{\sup}_{x\in\mathbb{R}}\frac{|b_{N}(t\,,x)|}{1+|x|}\leq L_{b}<\infty\quad\text{ and }\quad\adjustlimits{\sup}_{t>0}{\sup}_{x\in\mathbb{R}}\frac{|\sigma_{N}(t\,,x)|}{1+|x|}\leq L_{\sigma}<\infty,$
uniformly in $N>0$. Thanks to standard theory [Dalang, Walsh], the following SPDE has a predictable mild solution: For $(t\,,x)\in(0\,,\infty)\times\mathbb{R}$,
$\partial_{t}u_{N}(t\,,x)=\tfrac{1}{2}\partial^{2}_{x}u_{N}(t\,,x)+b_{N}(t\,,u_{N}(t\,,x))+\sigma_{N}(t\,,u_{N}(t\,,x))\dot{W}(t\,,x),$ (2.2)
subject to $u_{N}(0\,,x)=u_{0}(x)$. Moreover this solution is unique subject to the condition,
$\adjustlimits{\sup}_{t\in(0,T)}{\sup}_{x\in\mathbb{R}}\mathrm{E}\left(|u_{N}(t\,,x)|^{k}\right)<\infty\quad\text{for all $N,T>0$ and $k\geq 1$}.$ (2.3)
See Dalang and Sanz-Solé [bookDS, Theorem 4.2.1]. As usual, (2.2) is short-hand for the random integral equation,
$u_{N}(t\,,x)=(p_{t}*u_{0})(x)+\mathcal{I}^{N}_{b}(t\,,x)+\mathcal{I}^{N}_{\sigma}(t\,,x),$ (2.4)
where $\mathcal{I}^{N}_{b}$ and $\mathcal{I}^{N}_{\sigma}$ are the truncated analogous of the integrals in (1.4), where for every space-time function $\psi$,
$\mathcal{I}^{N}_{\psi}(t\,,x)=\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\psi_{N}(s\,,u_{N}(s\,,y))\,\mathrm{d}s\,\mathrm{d}y.$ (2.5)
The next result yields a moment estimate of the truncated solution which is similar to that in Theorem 6.3.2 of [bookDS] for more general SPDEs. We describe the details in order to provide the explicit constants, and parameter dependencies, of the bound.
Proposition 4.
If $L_{\sigma}>0$, then
$\adjustlimits{\sup}_{N>0}{\sup}_{x\in\mathbb{R}}\mathrm{E}\left(|u_{N}(t\,,x)|^{k}\right)\leq 4^{k}(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+1)^{k}\mathrm{e}^{128L_{\sigma}^{4}k^{3}t}.$
uniformly for all $t>0$ and $k\geq\max(2\,,L_{b}^{1/2}L_{\sigma}^{-2})$.
Proof.
Choose and fix $N,t>0$ and $x\in\mathbb{R}$. Owing to (2.4), we may write
$\|u_{N}(t\,,x)\|_{k}\leq\|u_{0}\|_{L^{\infty}(\mathbb{R})}+I_{1}+I_{2},$ (2.6)
where $I_{1}=\|\mathcal{I}^{N}_{b}(t\,,x)\|_{k}$ and $I_{2}=\|\mathcal{I}^{N}_{\sigma}(t\,,x)\|_{k}.$ We begin by estimating $I_{1}$ as follows. Thanks to (1.2) and the Minkowski inequality and the fact that $\|b_{N}(s\,,u_{N}(s\,,y))\|_{k}\leq L_{b}(1+\|u_{N}(s\,,y)\|_{k})$,
$I_{1}$ $\leq\int_{0}^{t}\mathrm{d}s\int_{-\infty}^{\infty}\mathrm{d}y\ p_{t-s}(y-x)\left\|b_{N}(s\,,u_{N}(s\,,y))\right\|_{k}$
$\leq L_{b}\left[t+\int_{0}^{t}\sup_{y\in\mathbb{R}}\|u_{N}(s\,,y)\|_{k}\,\mathrm{d}s\right].$
For every space-time random field $Z=\{Z(t\,,x);\,t\geq 0,x\in\mathbb{R}\}$ and for all real numbers $k\geq 1$ and $\beta>0$ define
$\mathcal{N}_{k,\beta}(Z)=\sup_{t\geq 0}\sup_{x\in\mathbb{R}}\mathrm{e}^{-\beta t}\|Z(t\,,x)\|_{k}.$ (2.7)
It follows that
$I_{1}\leq L_{b}\left[t+\mathcal{N}_{k,\beta}(u_{N})\int_{0}^{t}\mathrm{e}^{\beta s}\,\mathrm{d}s\right]\leq L_{b}\left[t+\frac{\mathrm{e}^{\beta t}}{\beta}\mathcal{N}_{k,\beta}(u_{N})\right].$
Since $t\exp(-\beta t)\leq(\mathrm{e}\beta)^{-1}<\beta^{-1}$, this leads to the following bound for $I_{1}$:
$I_{1}\leq\frac{L_{b}\mathrm{e}^{\beta t}}{\beta}\left[1+\mathcal{N}_{k,\beta}(u_{N})\right].$ (2.8)
We bound $I_{2}$ using the asymptotically optimal form of the Burkholder-Davis-Gundy inequality (see [minicourse]) as follows:
$I_{2}^{2}\leq 4k\int_{0}^{t}\mathrm{d}s\int_{-\infty}^{\infty}\mathrm{d}y\ \left[p_{t-s}(y-x)\right]^{2}\left\|\sigma_{N}(s\,,u_{N}(s\,,y))\right\|_{k}^{2}.$ (2.9)
Thanks to (1.2) and the fact that $\|\sigma_{N}(s\,,u_{N}(s\,,y))\|\leq 2L_{\sigma}^{2}(1+\|u_{N}(s\,,y)\|_{k}^{2})$,
$I_{2}^{2}$ $\leq 8kL_{\sigma}^{2}\int_{0}^{t}\mathrm{d}s\int_{-\infty}^{\infty}\mathrm{d}y\ \left[p_{t-s}(y-x)\right]^{2}\left(1+\|u_{N}(s\,,y)\|_{k}^{2}\right).$
Basic properties of the heat kernel imply that
$\|p_{r}\|_{L^{2}(\mathbb{R})}^{2}=(p_{r}*p_{r})(0)=p_{2r}(0)=\frac{1}{2\sqrt{\pi r}}\qquad\text{for every $r>0$}.$ (2.10)
Therefore,
$I_{2}^{2}$ $\leq\frac{4kL_{\sigma}^{2}}{\sqrt{\pi}}\int_{0}^{t}\frac{\mathrm{d}s}{\sqrt{t-s}}+\frac{4kL_{\sigma}^{2}}{\sqrt{\pi}}\int_{0}^{t}\sup_{y\in\mathbb{R}}\|u_{N}(s\,,y)\|_{k}^{2}\,\frac{\mathrm{d}s}{\sqrt{t-s}}$
$\leq\frac{4kL_{\sigma}^{2}}{\sqrt{\pi}}\int_{0}^{t}\frac{\mathrm{d}s}{\sqrt{s}}+\frac{4kL_{\sigma}^{2}\left[\mathcal{N}_{k,\beta}(u_{N})\right]^{2}\mathrm{e}^{2\beta t}}{\sqrt{\pi}}\int_{0}^{t}\frac{\mathrm{e}^{-2\beta(t-s)}}{\sqrt{t-s}}\,\mathrm{d}s.$
Since $\int_{0}^{t}s^{-1/2}\,\mathrm{d}s\leq\exp(2\beta t)\int_{0}^{\infty}s^{-1/2}\exp(-2\beta s)\,\mathrm{d}s=\sqrt{\pi/(2\beta)}\exp(2\beta t),$ we are led to the following:
$I_{2}^{2}\leq\frac{4k\mathrm{e}^{2\beta t}L_{\sigma}^{2}}{\sqrt{2\beta}}\left(1+\left[\mathcal{N}_{k,\beta}(u_{N})\right]^{2}\right).$
We prefer to simplify the preceding slightly more, using the inequality $\sqrt{l^{2}+n^{2}}\leq|l|+|n|$ – valid for all $l,n\in\mathbb{R}$ – as follows:
$I_{2}\leq\frac{2\sqrt{k}\mathrm{e}^{\beta t}L_{\sigma}}{(2\beta)^{1/4}}\left(1+\mathcal{N}_{k,\beta}(u_{N})\right).$ (2.11)
We can combine (2.8) and (2.11) to find that
$\|u_{N}(t\,,x)\|_{k}$ $\leq\|u_{0}\|_{L^{\infty}(\mathbb{R})}+\frac{L_{b}\mathrm{e}^{\beta t}}{\beta}\left[1+\mathcal{N}_{k,\beta}(u_{N})\right]+\frac{2\sqrt{k}\mathrm{e}^{\beta t}L_{\sigma}}{(2\beta)^{1/4}}\left(1+\mathcal{N}_{k,\beta}(u_{N})\right)$
$\leq\|u_{0}\|_{L^{\infty}(\mathbb{R})}+\mathrm{e}^{\beta t}\left(\frac{L_{b}}{\beta}+\frac{2\sqrt{k}L_{\sigma}}{(2\beta)^{1/4}}\right)\left(1+\mathcal{N}_{k,\beta}(u_{N})\right).$
This implies that
$\mathcal{N}_{k,\beta}(u_{N})\leq\|u_{0}\|_{L^{\infty}(\mathbb{R})}+\left(\frac{L_{b}}{\beta}+\frac{2\sqrt{k}L_{\sigma}}{(2\beta)^{1/4}}\right)\left(1+\mathcal{N}_{k,\beta}(u_{N})\right).$
Set $\beta=128k^{2}L_{\sigma}^{4}$ to see that $\beta^{-1}L_{b}+(2\beta)^{-1/4}2\sqrt{k}L_{\sigma}\leq\frac{3}{4}$ for this particular choice of $\beta$. Solve for $\mathcal{N}_{k,128k^{2}L_{\sigma}^{4}}(u_{N})$ in order to find that $\mathcal{N}_{k,128k^{2}L_{\sigma}^{4}}(u_{N})\leq 4(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+1).$ This is another way to state the proposition. ∎
Proposition 4 omits the case that $\sigma$ is a constant. The following takes care of that case, and at the same time improves on Proposition 4 when $\sigma$ is bounded but $\text{Lip}(\sigma)>0$.
Proposition 5.
If $\sigma:(0\,,\infty)\times\mathbb{R}\to\mathbb{R}$ is bounded, then
$\adjustlimits{\sup}_{N>0}{\sup}_{x\in\mathbb{R}}\mathrm{E}\left(|u_{N}(t\,,x)|^{k}\right)\leq 4^{k}\mathrm{e}^{2L_{b}kt}\left(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}t^{1/4}+1\right)^{k}k^{k/2},$
uniformly for all $t>0$ and $k\geq 2$.
Proof.
We modify the proof of Proposition 4 by first observing that (2.6) remains valid, and so does (2.8). We start with (2.9) and estimate $I_{2}$, using (2.10), simply as follows:
$I_{2}^{2}$ $\leq 4k\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}^{2}\int_{0}^{t}\|p_{r}\|_{L^{2}(\mathbb{R})}^{2}\,\mathrm{d}r=\frac{2k}{\sqrt{\pi}}\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}^{2}\int_{0}^{t}\frac{\mathrm{d}r}{\sqrt{r}}$
$=4k\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}^{2}\sqrt{t/\pi}\leq 4k\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}^{2}\sqrt{t}.$
This yields
$\|u_{N}(t\,,x)\|_{k}\leq\|u_{0}\|_{L^{\infty}(\mathbb{R})}+\frac{L_{b}\exp(\beta t)}{\beta}\left[1+\mathcal{N}_{k,\beta}(u_{N})\right]+2\sqrt{k}\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}t^{1/4}.$
Since the right-hand side does not depend on $(t\,,x)$, we divide by $\exp(\beta t)$ and optimize over $(t\,,x)$ to find that
$\mathcal{N}_{k,\beta}(u_{N})\leq\|u_{0}\|_{L^{\infty}(\mathbb{R})}+2\sqrt{k}\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}t^{1/4}+\frac{L_{b}}{\beta}\left[1+\mathcal{N}_{k,\beta}(u_{N})\right],$
uniformly for all real numbers $k\geq 2$, $N,\beta>0$. Set $\beta=2L_{b}$ and solve for $\mathcal{N}_{k,2L_{b}}(u_{N})$ to find that
$\mathcal{N}_{k,2L_{b}}(u_{N})$ $\leq 2\|u_{0}\|_{L^{\infty}(\mathbb{R})}+4\sqrt{k}\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}t^{1/4}+1$
$\leq 4\left(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}t^{1/4}+1\right)\sqrt{k},$
which is another way to state the result. ∎
The next results shows some tail estimates of the truncated solution.
Proposition 6.
If $L_{\sigma}>0$, then
$\mathrm{P}\left\{|u_{N+1}(t\,,x)|\geq\mathrm{e}^{N}\right\}\leq\exp\left(-\frac{N^{3/2}}{(512t)^{1/2}L_{\sigma}^{2}}\right),$
uniformly for all $t>0$, $x\in\mathbb{R}$, and
$N\geq 4\log(4(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+1))\vee 256t\max\left(4L_{\sigma}^{4},L_{b}\right).$
If $\sigma\in L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})$, then
$\mathrm{P}\left\{|u_{N+1}(t\,,x)|\geq\mathrm{e}^{N}\right\}\leq\exp\left(-\frac{\mathrm{e}^{2N-4L_{b}t}}{32\mathrm{e}\left(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}t^{1/4}+1\right)^{2}}\right),$
uniformly for all $t>0$, $x\in\mathbb{R}$, and
$N\geq\frac{1}{2}\log 32+2L_{b}t+\frac{1}{2}+\log\left(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}t^{1/4}+1\right).$
Proof.
Let us first consider the case that $L_{\sigma}>0$. Proposition 4 and Chebyshev’s inequality together ensure that
$\begin{split}\mathrm{P}\left\{|u_{N+1}(t\,,x)|\geq\mathrm{e}^{N}\right\}&\leq\mathrm{e}^{-kN}\mathrm{E}\left(|u_{N+1}(t\,,x)|^{k}\right)\\ &\leq\left(4(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+1)\right)^{k}\mathrm{e}^{-kN+128L_{\sigma}^{4}k^{3}t},\end{split}$
uniformly for all real numbers $N,t>0$, $x\in\mathbb{R}$, and all real numbers $k\geq\max(2\,,L_{b}^{1/2}L_{\sigma}^{-2})$. Set $k=\sqrt{AN}$ – where the value of $A>0$ is to be determined – and $C=4(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+1)$ in order to see that
$\mathrm{P}\left\{|u_{N+1}(t\,,x)|\geq\mathrm{e}^{N}\right\}\leq\exp\left(-\left\{\sqrt{A}-128L_{\sigma}^{4}A^{3/2}t-\frac{\sqrt{A}\log C}{N}\right\}N^{3/2}\right).$
We use this with $A=(256L_{\sigma}^{4}t)^{-1}$, and observe that $k\geq\max(2\,,\sqrt{L_{b}}L_{\sigma}^{-2})$ iff $N\geq 256t\max(4L_{\sigma}^{4},L_{b})$. It follows that
$\mathrm{P}\left\{|u_{N+1}(t\,,x)|\geq\mathrm{e}^{N}\right\}$ $\leq\exp\left(-\left\{\frac{128L_{\sigma}^{4}t}{[256L_{\sigma}^{4}t]^{3/2}}-\frac{\log C}{N\sqrt{256L_{\sigma}^{4}t}}\right\}N^{3/2}\right)$
$=\exp\left(-\left\{\frac{1}{2}-\frac{\log C}{N}\right\}\frac{N^{3/2}}{(256t)^{1/2}L_{\sigma}^{2}}\right),$
which has the desired outcome, when $L_{\sigma}>0$, provided additionally that $N\geq 4\log C$. The case $\sigma\in L^{\infty}(\mathbb{R})$ is proved similarly but rests on Proposition 5 instead of 4, viz.,
$\mathrm{P}\left\{|u_{N+1}(t\,,x)|\geq\mathrm{e}^{N}\right\}\leq\mathrm{e}^{-kN}\mathrm{E}\left(|u_{N+1}(t\,,x)|^{k}\right)\leq C^{k}\mathrm{e}^{-kN}k^{k/2},$
valid uniformly for all $N,t>0$, $x\in\mathbb{R}$, $k\geq 2$, where
$C=C(t\,,u_{0}\,,L_{b})=4\mathrm{e}^{2L_{b}t}\left(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}t^{1/4}+1\right).$
We apply the preceding with the particular choice $k=C^{-2}\exp\{2N-1\}$ and compute a bit in order to finish. ∎
Finally, the following real-variable lemma will be a key result for the proof of existence of the solution to (1.1).
Lemma 7.
Choose and fix two strictly positive real numbers $\alpha$ and $\gamma$ that satisfy $\gamma>\alpha^{-1}$. Then, $\sum_{N=1}^{\infty}a_{N}^{-\alpha}\exp(-N^{\gamma}/a_{N})<\infty$ for every non-decreasing sequence $0<a_{1}\leq a_{2}\leq a_{3}\leq\cdots$ of positive numbers.
Proof.
Choose and fix a number $\rho\in(\alpha^{-1}\!,\gamma)$ and divide the sum according to whether or not $a_{N}\leq N^{\rho}$. The lemma is true because $\sum_{a_{N}>N^{-\rho}}\leq\sum_{a_{N}>N^{\rho}}a_{N}^{-\alpha}\leq\sum_{a_{N}>N^{\rho}}N^{-\alpha\rho}$ and $\sum_{a_{N}\leq N^{-\rho}}a_{N}^{-\alpha}\exp(-N^{\gamma}/a_{N})\leq a_{1}^{-\alpha}\sum_{a_{N}\leq N^{\rho}}\exp(-N^{\gamma-\rho})$.
3. Proof of Theorem 3
Proof of existence.
We will prove the theorem in the case that $L_{\sigma}>0$. Small adjustments to proof yield the result also in the remaining case that $\sigma$ is bounded.
Throughout the proof we choose and fix a non-random number $T>0$, and then strive to prove that (1.1) has a mild solution up to time $T$. Since $T$ is arbitrary, this proves the theorem. With this number in mind, we adjust (2.7) slightly and define
$\mathcal{N}_{k,\beta,T}(Z)=\sup_{t\in(0,T)}\sup_{x\in\mathbb{R}}\mathrm{e}^{-\beta t}\|Z(t\,,x)\|_{k},$
for every $\beta>0$, $k\geq 1$, and all space-time random fields $Z$.
Note that $N\mapsto\text{\rm Lip}_{N}(b)$ and $N\mapsto\text{\rm Lip}_{N}(\sigma)$ are nondecreasing and $b$ and $\sigma$ are globally Lipschitz when $\lim_{N\to\infty}\text{\rm Lip}_{N}(b)<\infty$ and $\lim_{N\to\infty}\text{\rm Lip}_{N}(\sigma)<\infty$. Therefore, we assume without loss of generality that
$\lim_{N\to\infty}L_{N,b}=\infty\quad\text{and}\quad\lim_{N\to\infty}L_{N,\sigma}=\infty,$ (3.1)
where $L_{N,b}$ and $L_{N,\sigma}$ were defined in (2.1).
Thanks to (2.4) we can write
$\left\|u_{N+1}(t\,,x)-u_{N}(t\,,x)\right\|_{k}\leq I_{1}+I_{2},$ (3.2)
where
$I_{1}$ $=\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\left\|b_{N+1}(s\,,u_{N+1}(s\,,y))-b_{N}(s\,,u_{N}(s\,,y))\right\|_{k}\mathrm{d}s\,\mathrm{d}y,$
$I_{2}$ $=\left\|\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\left[\sigma_{N+1}(s\,,u_{N+1}(s\,,y))-\sigma_{N}(s\,,u_{N}(s\,,y))\right]W(\mathrm{d}s\,\mathrm{d}y)\right\|_{k}.$
Recall (2.7) and for every $N,s>0$ and $y\in\mathbb{R}$ consider the event
$G_{N+1}(s\,,y)=\left\{\omega\in\Omega:\,|u_{N+1}(s\,,y)|(\omega)\leq\mathrm{e}^{N}\right\}.$
On one hand,
$\left\|\left[b_{N+1}(s\,,u_{N+1}(s\,,y))-b_{N}(s\,,u_{N}(s\,,y))\right]\mathbf{1}_{G_{N+1}(s\,,y)}\right\|_{k}$
$\leq\left\|b_{N}(s\,,u_{N+1}(s\,,y))-b_{N}(s\,,u_{N}(s\,,y))\right\|_{k}$
$\leq L_{N,b}\left\|u_{N+1}(s\,,y)-u_{N}(s\,,y)\right\|_{k}{\color[rgb]{0,0,0}\leq L_{N,b}\mathrm{e}^{\beta s}\mathcal{N}_{k,\beta,T}(u_{N+1}-u_{N}),}$
for all $\beta,N>0$, $s\in(0\,,T)$, and $y\in\mathbb{R}$. On the other hand,
$\left\|\left[b_{N+1}(s\,,u_{N+1}(s\,,y))-b_{N}(s\,,u_{N}(s\,,y))\right]\mathbf{1}_{\Omega\setminus G_{N+1}(s\,,y)}\right\|_{k}$
$\leq\left\|b_{N+1}(s\,,u_{N+1}(s\,,y))\mathbf{1}_{\Omega\setminus G_{N+1}(s\,,y)}\right\|_{k}+\left\|b_{N}(s\,,u_{N}(s\,,y))\mathbf{1}_{\Omega\setminus G_{N+1}(s\,,y)}\right\|_{k}$
$\leq\left[\left\|b_{N+1}(s\,,u_{N+1}(s\,,y))\right\|_{2k}+\left\|b_{N}(s\,,u_{N}(s\,,y))\right\|_{2k}\right]\left[1-\mathrm{P}\left(G_{N+1}(s\,,y)\right)\right]^{1/(2k)}.$
We have used the following variation of the Cauchy–Schwarz inequality in the last line: $\|X\mathbf{1}_{F}\|_{k}\leq\|X\|_{2k}[\mathrm{P}(F)]^{1/(2k)}$ for all $X\in L^{k}(\Omega)$ and all events $F\subset\Omega$. Set $C=4(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+1).$ Thanks to Proposition 4,
$\left\|b_{N+1}(s\,,u_{N+1}(s\,,y))\right\|_{2k}+\left\|b_{N}(s\,,u_{N}(s\,,y))\right\|_{2k}$
$\leq L_{b}\left[\|u_{N+1}(s\,,y)\|_{2k}+\|u_{N}(s\,,y)\|_{2k}\right]\leq{\color[rgb]{0,0,0}2CL_{b}}\mathrm{e}^{512L_{\sigma}^{4}k^{2}s},$
uniformly for all $N,s>0$, $y\in\mathbb{R}$, and $k\geq\max(1\,,\frac{1}{2}L_{b}^{1/2}L_{\sigma}^{-2})$. This yields the following inequality:
$\left\|\left[b_{N+1}(s\,,u_{N+1}(s\,,y))-b_{N}(s\,,u_{N}(s\,,y))\right]\mathbf{1}_{\Omega\setminus G_{N+1}(s\,,y)}\right\|_{k}$
$\qquad\leq{\color[rgb]{0,0,0}2CL_{b}}\mathrm{e}^{512L_{\sigma}^{4}k^{2}s}\left[\mathrm{P}\left\{|u_{N+1}(s\,,y)|\geq\mathrm{e}^{N}\right\}\right]^{1/(2k)},$
valid uniformly for all $N,s>0$, $y\in\mathbb{R}$, and $k\geq\max(1\,,\frac{1}{2}L_{b}^{1/2}L_{\sigma}^{-2})$. Therefore, Proposition 6 yields
$\left\|\left[b_{N+1}(s\,,u_{N+1}(s\,,y))-b_{N}(s\,,u_{N}(s\,,y))\right]\mathbf{1}_{\Omega\setminus G_{N+1}(s\,,y)}\right\|_{k}$
$\qquad\leq 2L_{b}C\mathrm{e}^{512L_{\sigma}^{4}k^{2}s}\exp\left(-\frac{N^{3/2}}{2k(512s)^{1/2}L_{\sigma}^{2}}\right),$
valid uniformly for all $s\in(0\,,T)$, $y\in\mathbb{R}$, $N\geq 4\log C\vee 256T\max\left(4L_{\sigma}^{4},L_{b}\right)$, and $k\geq\max(1\,,\frac{1}{2}L_{b}^{1/2}L_{\sigma}^{-2})$. Thus, we find that
$I_{1}$ $\leq L_{N,b}\mathrm{e}^{\beta t}\mathcal{N}_{k,\beta,T}(u_{N+1}-u_{N})\int_{(0,t)\times\mathbb{R}}\mathrm{e}^{-\beta(t-s)}p_{t-s}(y-x)\,\mathrm{d}s\,\mathrm{d}y$
$\quad+2L_{b}C\int_{(0,t)\times\mathbb{R}}\mathrm{e}^{512L_{\sigma}^{4}k^{2}s}\exp\left(-\frac{N^{3/2}}{2k(512s)^{1/2}L_{\sigma}^{2}}\right)p_{t-s}(y-x)\,\mathrm{d}s\,\mathrm{d}y$ (3.3)
$\leq\frac{L_{N,b}\mathrm{e}^{\beta t}}{\beta}\,\mathcal{N}_{k,\beta,T}(u_{N+1}-u_{N})+\frac{L_{b}C}{256L_{\sigma}^{4}k^{2}}\,\exp\left(512L_{\sigma}^{4}k^{2}t-\frac{N^{3/2}}{2k(512t)^{1/2}L_{\sigma}^{2}}\right),$
uniformly for all $t\in(0\,,T)$ and $x\in\mathbb{R}$, provided that $N,k\geq c$, for a complicated looking but otherwise unimportant number $c=c(\|u_{0}\|_{L^{\infty}(\mathbb{R})}\,,L_{\sigma}\,,T\,,L_{b})>0$.
Next we study the quantity $I_{2}$. Let us appeal to the Burkholder-Davis-Gundy inequality (see [minicourse]) to find that
$I_{2}^{2}$ $\leq 4k\int_{0}^{t}\mathrm{d}s\int_{-\infty}^{\infty}\mathrm{d}y\ [p_{t-s}(y-x)]^{2}\left\|\sigma_{N+1}(s\,,u_{N+1}(s\,,y))-\sigma_{N}(s\,,u_{N}(s\,,y))\right\|_{k}^{2}.$
As in the above truncation, we may write
$\left\|\left[\sigma_{N+1}(s\,,u_{N+1}(s\,,y))-\sigma_{N}(s\,,u_{N}(s\,,y))\right]\mathbf{1}_{G_{N+1}(s\,,y)}\right\|_{k}$
$\leq L_{N,\sigma}\mathrm{e}^{\beta s}\mathcal{N}_{k,\beta,T}(u_{N+1}-u_{N})\quad\text{ for all $\beta,N>0$, $s\in(0\,,T)$, and $y\in\mathbb{R}$.}$
Moreover,
$\left\|\left[\sigma_{N+1}(s\,,u_{N+1}(s\,,y))-\sigma_{N}(s\,,u_{N}(s\,,y))\right]\mathbf{1}_{\Omega\setminus G_{N+1}(s\,,y)}\right\|_{k}$
$\hskip 7.22743pt\leq\left[\left\|\sigma_{N+1}(s\,,u_{N+1}(s\,,y))\right\|_{2k}+\left\|\sigma_{N}(s\,,u_{N}(s\,,y))\right\|_{2k}\right]\left[1-\mathrm{P}\left(G_{N+1}(s\,,y)\right)\right]^{1/(2k)}.$
Thanks to Proposition 4,
$\left\|\sigma_{N+1}(s\,,u_{N+1}(s\,,y))\right\|_{2k}+\left\|\sigma_{N}(s\,,u_{N}(s\,,y))\right\|_{2k}$
$\hskip 72.26999pt\leq L_{\sigma}\left[\|u_{N+1}(s\,,y)\|_{2k}+\|u_{N}(s\,,y)\|_{2k}\right]\leq 2L_{\sigma}C\mathrm{e}^{512L_{\sigma}^{4}k^{2}s},$
uniformly for all $N,s>0$, $y\in\mathbb{R}$, and $k\geq\max(1\,,\frac{1}{2}L_{b}^{1/2}L_{\sigma}^{-2})$. Therefore, Proposition 6 yields
$\left\|\left[\sigma_{N+1}(s\,,u_{N+1}(s\,,y))-\sigma_{N}(s\,,u_{N}(s\,,y))\right]\mathbf{1}_{\Omega\setminus G_{N+1}(s\,,y)}\right\|_{k}$
$\hskip 122.85876pt\leq 2L_{\sigma}C\mathrm{e}^{512L_{\sigma}^{4}k^{2}s}\exp\left(-\frac{N^{3/2}}{2k(512s)^{1/2}L_{\sigma}^{2}}\right),$
uniformly for all $s\in(0\,,T)$, $y\in\mathbb{R}$, $N\geq 4\log C\vee 256T\max\left(4L_{\sigma}^{4},L_{b}\right)$, and $k\geq\max(1\,,\frac{1}{2}L_{b}^{1/2}L_{\sigma}^{-2})$. Thus, we find that
$I_{2}^{2}\leq 8kL_{N,\sigma}^{2}\mathrm{e}^{2\beta t}\left[\mathcal{N}_{k,\beta,T}(u_{N+1}-u_{N})\right]^{2}\int_{0}^{t}\mathrm{d}s\int_{-\infty}^{\infty}\mathrm{d}y\ \mathrm{e}^{-2\beta(t-s)}[p_{t-s}(y-x)]^{2}$
$\quad+32kL_{b}^{2}C^{2}\int_{(0,t)\times\mathbb{R}}\mathrm{e}^{1024L_{\sigma}^{4}k^{2}s}\exp\left(-\frac{N^{3/2}}{k(512s)^{1/2}L_{\sigma}^{2}}\right)[p_{t-s}(y-x)]^{2}\,\mathrm{d}s\,\mathrm{d}y$
$\leq\frac{4kL_{N,\sigma}^{2}\mathrm{e}^{2\beta t}}{\sqrt{\beta}}\left[\mathcal{N}_{k,\beta,T}(u_{N+1}-u_{N})\right]^{2}{\color[rgb]{0,0,0}+32kL_{\sigma}^{2}C^{2}\mathrm{e}^{1024L_{\sigma}^{4}k^{2}t}\exp\left(-\frac{N^{3/2}}{k(512t)^{1/2}L_{\sigma}^{2}}\right)\sqrt{t},}$
for every choice of $\beta>0$, $t\in(0\,,T)$, and $x\in\mathbb{R}$, provided that $N,k\geq c$. Combine this last inequality with (3.3) and (3.2) in order to see that
$\|u_{N+1}(t\,,x)-u_{N}(t\,,x)\|_{k}\leq\mathrm{e}^{\beta t}\left[\frac{L_{N,b}}{\beta}+\frac{2\sqrt{k}L_{N,\sigma}}{\beta^{1/4}}\right]\mathcal{N}_{k,\beta,T}(u_{N+1}-u_{N})$
$\qquad+\frac{L_{b}C}{256L_{\sigma}^{4}k^{2}}\,\exp\left(512L_{\sigma}^{4}k^{2}t-\frac{N^{3/2}}{2k(512t)^{1/2}L_{\sigma}^{2}}\right)$
$\qquad+\sqrt{32k}L_{\sigma}C\mathrm{e}^{512L_{\sigma}^{4}k^{2}t}\exp\left(-\frac{N^{3/2}}{2k(512t)^{1/2}L_{\sigma}^{2}}\right)t^{1/4}.$
as long as $N,k\geq c$. Set $\beta=512L_{N,\sigma}^{4}k^{6}$, divide both sides by $\exp(\beta t)$ and optimize over $(t\,,x)$ in order to see that
$\mathcal{N}_{k,512L_{N,\sigma}^{4}k^{6},T}(u_{N+1}-u_{N})\leq\left[\frac{L_{N,b}}{4^{4}L_{N,\sigma}^{4}k^{6}}+\frac{1}{4k^{11/2}}\right]\mathcal{N}_{k,512L_{\sigma}^{4}k^{6},T}(u_{N+1}-u_{N})$
$\qquad\qquad+\frac{L_{b}C}{256L_{\sigma}^{4}k^{2}}\,\exp\left(512t(L_{\sigma}^{4}k^{2}-L_{N,\sigma}^{4}k^{6})-\frac{N^{3/2}}{2k(512T)^{1/2}L_{\sigma}^{2}}\right)$ (3.4)
$\qquad\qquad+\sqrt{32k}t^{1/4}L_{\sigma}C\mathrm{e}^{512t(L_{\sigma}^{4}k^{2}-L_{N,\sigma}^{4}k^{6})}\exp\left(-\frac{N^{3/2}}{2k(512T)^{1/2}L_{\sigma}^{2}}\right).$
Set $k=\sqrt{L_{N,b}}$ and take $N$ large enough such that $\alpha_{0}:=(4^{4}L_{N,\sigma}^{4}L_{N,b}^{2})^{-1}+(4L_{N,b}^{11/4})^{-1}<1$ and $L_{\sigma}^{4}<L_{N,\sigma}^{4}$. In particular, if $k^{4}\geq 2$, we observe that $\exp\{512t(L_{\sigma}^{4}k^{2}-L_{N,\sigma}^{4}k^{6})\}\leq\exp\{-512tL_{N,\sigma}^{4}k^{6}(1-\frac{1}{k^{4}})\}\leq\exp\{-256tk^{6}L_{N,\sigma}^{4}\}$. Since $\exp(-x)\leq x^{-1/4}$ when $x>0$, we find that
$\adjustlimits{\sup}_{t\in(0,T)}{\sup}_{x\in\mathbb{R}}\|u_{N+1}(t\,,x)-u_{N}(t\,,x)\|_{\sqrt{L_{N,b}}}$
$\hskip 36.135pt\leq\frac{1}{(1-\alpha_{0})}\frac{L_{b}C}{256L_{\sigma}^{4}L_{N,b}}\,\exp\left(-\frac{N^{3/2}}{2\sqrt{L_{N,b}}(512T)^{1/2}L_{\sigma}^{2}}\right)$
$\hskip 50.58878pt+\frac{1}{(1-\alpha_{0})}\frac{2C}{L_{N,b}}\exp\left(-\frac{N^{3/2}}{2\sqrt{L_{N,b}}(512T)^{1/2}L_{\sigma}^{2}}\right),$
provided only that $N,k\geq c$ [^1] It is possible to use a priori estimates to show that the above is the only possible subsolution to (3.4); in other words, that $\mathcal{N}_{k,512L_{N,\sigma}^{4}k^{2},T}(u_{N+1}-u_{N})$ is not infinite. The method uses Picard iteration and follows along similar lines that were developed in Foondun and Khoshnevisan [FK1]. We skip the several lines of additional technical details. Thanks to (3.1), the above bounds $\|u_{N+1}(t\,,x)-u_{N}(t\,,x)\|_{2}$ for all large $N$, and we find that
$\adjustlimits{\sup}_{t\in(0,T)}{\sup}_{x\in\mathbb{R}}\|u_{N+1}(t\,,x)-u_{N}(t\,,x)\|_{2}\lesssim L_{N,b}^{-1}\exp\left(-\frac{N^{3/2}}{2(512T)^{1/2}L_{\sigma}^{2}\sqrt{L_{N,b}}}\right),$
uniformly for all $N\geq c^{\prime}=c^{\prime}(\|u_{0}\|_{L^{\infty}(\mathbb{R})}\,,L_{\sigma}\,,L_{b}\,,T)>0$ for a complicated looking but otherwise unimportant number $c^{\prime}$. We appeal to Lemma 7 [with $a_{N}=2(512T)^{1/2}L_{\sigma}^{2}\sqrt{L_{N,b}}$, $\gamma=3/2$ and $\alpha=2$] to see that
$\sum_{\begin{subarray}{c}N\in\mathbb{N}\cap[c^{\prime},\infty)\end{subarray}}\adjustlimits{\sup}_{t\in(0,T)}{\sup}_{x\in\mathbb{R}}\|u_{N+1}(t\,,x)-u_{N}(t\,,x)\|_{2}<\infty.$
Thus, we may deduce from (2.3) and (3.1) that the above continues to hold if we replace the sum with the sum over all $N\in\mathbb{N}$. This proves that
$u(t\,,x)=\lim_{N\to\infty}u_{N}(t\,,x)\quad\text{exists in $L^{2}(\Omega)$},$ (3.5)
and the rate of convergence does not depend on $t\in(0\,,T)$ nor on $x\in\mathbb{R}$. The random field $u$ is predictable since it is a limit of predictable random fields. Therefore it remains to prove that, for every $t\in(0\,,T)$ and $x\in\mathbb{R}$,
$\lim_{N\to\infty}\mathcal{I}^{N}_{b}(t\,,x)=\mathcal{I}_{b}(t\,,x)\quad\text{and}\quad\lim_{N\to\infty}\mathcal{I}^{N}_{\sigma}(t\,,x)=\mathcal{I}_{\sigma}(t\,,x),$ (3.6)
where both of the limits hold in $L^{2}(\Omega)$, and the random fields $\mathcal{I}_{b}$, $\mathcal{I}_{\sigma}$, $\mathcal{I}^{N}_{b}$, and $\mathcal{I}^{N}_{\sigma}$ were defined respectively in (1.4) and (2.5). Thanks to (3.5) and (2.4), this proves that $u$ is a mild solution to (1.1).
Observe that $|b_{N}(t\,,z)-b(t\,,z)|\leq\mathbf{1}_{\mathbb{R}\setminus[-\exp(N),\exp(N)]}(z)(|b(t\,,z)|+|b_{N}(t\,,z)|)\leq 2L_{b}(1+|z|)\mathbf{1}_{\mathbb{R}\setminus[-\exp(N),\exp(N)]}(z)$ for all $z\in\mathbb{R}$ and $t,N>0$. This yields
$\left\|\mathcal{I}^{N}_{b}(t\,,x)-\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)b(s\,,u_{N}(s\,,y))\,\mathrm{d}s\,\mathrm{d}y\right\|_{2}$
$\leq\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\left\|b_{N}(s\,,u_{N}(s\,,y))-b(s\,,u_{N}(s\,,y))\right\|_{2}\,\mathrm{d}s\,\mathrm{d}y$
$\leq 2L_{b}\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\left(\sqrt{\mathrm{E}\left(1+|u_{N}(s\,,y)|^{2};|u_{N}(s\,,y)|>\mathrm{e}^{N}\right)}\right)\,\mathrm{d}s\,\mathrm{d}y.$
If $X\geq 0$ is a random variable and $A>0$ is a constant, then we apply the Cauchy-Schwarz inequality and Chebyshev’s inequality back to back in order to find that $\mathrm{E}(X^{2};X>A)\leq\sqrt{\mathrm{E}(X^{4})\mathrm{P}\{|X|>A\}}\leq A^{-2}\mathrm{E}(X^{4}),$ Therefore, Propositions 4 and 5 assure us that
$\lim_{N\to\infty}\adjustlimits{\sup}_{s\in(0,T)}{\sup}_{y\in\mathbb{R}}\mathrm{E}\left(1+|u_{N}(s\,,y)|^{2};|u_{N}(s\,,y)|>\mathrm{e}^{N}\right)=0,$
and hence $\mathcal{I}^{N}_{b}(t\,,x)-\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)b(s\,,u_{N}(s\,,y))\,\mathrm{d}s\,\mathrm{d}y\to 0$ as $N\to\infty$, where the convergence takes place in $L^{2}(\Omega)$. Therefore, the first assertion of (3.6) would follow once we can show that
$\lim_{N\to\infty}\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)b(s\,,u_{N}(s\,,y))\,\mathrm{d}s\,\mathrm{d}y=\mathcal{I}_{b}(t\,,x),$ (3.7)
where the convergence takes place in $L^{2}(\Omega)$. Since $b$ has at-most linear growth and $\|b(s\,,u_{N}(s\,,y))-b(s\,,u(s\,,y))\|_{2}\leq\|b(s\,,u_{N}(s\,,y))\|_{2}+\|b(s\,,u(s\,,y))\|_{2},$ Propositions 4 and 5 ensure that $\|b(s\,,u_{N}(s\,,y))-b(s\,,u(s\,,y))\|_{2}$ is bounded uniformly in $s\in(0\,,T)$, $N>0$, and $y\in\mathbb{R}$. Thanks to (3.5), uniform integrability, and the continuity of $b$,
$\lim_{N\to\infty}b(s\,,u_{N}(s\,,y))=b(s\,,u(s\,,y))\quad\text{in $L^{2}(\Omega)$, for every $s>0$ and $y\in\mathbb{R}^{d}$. }$
Therefore, the dominated convergence theorem yields
$\lim_{N\to\infty}\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\left\|b(s\,,u_{N}(s\,,y))-b(s\,,u(s\,,y))\right\|_{2}\,\mathrm{d}s\,\mathrm{d}y=0,$
for all $t\in(0\,,T)$ and $x\in\mathbb{R}$. The triangle inequality now yields (3.7) and hence the first assertion of (3.6). Similarly,
$|\sigma_{N}(t\,,z)-\sigma(t\,,z)|\leq 2L_{\sigma}(1+|z|)\mathbf{1}_{\mathbb{R}\setminus[-\exp(N),\exp(N)]}(z)$
for all $z\in\mathbb{R}$ and $t,N>0$. Then, by the $L^{2}(\Omega)$-isometry of stochastic integrals,
$\left\|\mathcal{I}^{N}_{\sigma}(t\,,x)-\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\sigma(s\,,u_{N}(s\,,y))\,W(\mathrm{d}s\,\mathrm{d}y)\right\|_{2}^{2}$
$\leq\int_{(0,t)\times\mathbb{R}}[p_{t-s}(y-x)]^{2}\left\|\sigma_{N}(s\,,u_{N}(s\,,y))-\sigma(s\,,u_{N}(s\,,y))\right\|_{2}^{2}\,\mathrm{d}s\,\mathrm{d}y$
$\leq 4L_{\sigma}^{2}\int_{(0,t)\times\mathbb{R}}[p_{t-s}(y-x)]^{2}\mathrm{E}\left(1+|u_{N}(s\,,y)|^{2};|u_{N}(s\,,y)|>\mathrm{e}^{N}\right)\,\mathrm{d}s\,\mathrm{d}y.$
As before, Propositions 4 and 5 assure us that
$\lim_{N\to\infty}\adjustlimits{\sup}_{s\in(0,T)}{\sup}_{y\in\mathbb{R}}\mathrm{E}\left(1+|u_{N}(s\,,y)|^{2};|u_{N}(s\,,y)|>\mathrm{e}^{N}\right)=0,$
and hence $\mathcal{I}^{N}_{\sigma}(t\,,x)-\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\sigma(s\,,u_{N}(s\,,y))\,W(\mathrm{d}s\,\mathrm{d}y)\to 0$ in $L^{2}(\Omega)$ as $N\to\infty$. Therefore, we are left to show that
$\lim_{N\to\infty}\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\sigma(s\,,u_{N}(s\,,y))\,W(\mathrm{d}s\,\mathrm{d}y)=\mathcal{I}_{\sigma}(t\,,x),$
where the convergence takes place in $L^{2}(\Omega)$. Since $\sigma$ has at-most linear growth and
$\|\sigma(s\,,u_{N}(s\,,y))-\sigma(s\,,u(s\,,y))\|_{2}\leq\|\sigma(s\,,u_{N}(s\,,y))\|_{2}+\|\sigma(s\,,u(s\,,y))\|_{2},$
Propositions 4 and 5 ensure that $\|\sigma(s\,,u_{N}(s\,,y))-\sigma(s\,,u(s\,,y))\|_{2}$ is bounded uniformly in $s\in(0\,,T)$, $N>0$, and $y\in\mathbb{R}$. Thanks to (3.5), uniform integrability, and the continuity of $\sigma$, $\lim_{N\to\infty}\sigma(s\,,u_{N}(s\,,y))=\sigma(s\,,u(s\,,y))$ in $L^{2}(\Omega)$, for every $s>0$ and $y\in\mathbb{R}^{d}$. Therefore,
$\left\|\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\sigma(s\,,u_{N}(s\,,y))\,W(\mathrm{d}s\,\mathrm{d}y)-\mathcal{I}_{\sigma}(t\,,x)\right\|_{2}^{2}$
$\leq\int_{(0,t)\times\mathbb{R}}[p_{t-s}(y-x)]^{2}\left\|\sigma(s\,,u_{N}(s\,,y))-\sigma(s\,,u(s\,,y))\right\|_{2}^{2}\,\mathrm{d}s\,\mathrm{d}y,$
which converges to 0 as $N\rightarrow\infty$ by dominated convergence, for all $t\in(0\,,T)$ and $x\in\mathbb{R}$. The triangle inequality now yields (3.7). This verifies the second assertion of (3.6) and completes the proof. ∎
Proof of uniqueness.
As in the proof of existence, we only give the details of the proof in the case that $L_{\sigma}>0$, since the remaining case is similar.
Let $u,v$ two solutions to (1.1) with same initial condition $u_{0}$ satisfying Assumption 1. Recall that $u$ and $v$ satisfy the mild formulation (1.3) with $\sigma$ and $b$ satisfying Assumption 2. Then, using Burkholder-Davis-Gundy inequality (see [minicourse]), we find that $\|u(t\,,x)-v(t\,,x)\|_{k}\leq I_{1}+I_{2}+I_{3}+I_{4}$ for all $t>0$, $x\in\mathbb{R}$, and $k\geq 1$, where
$I_{1}$ $=\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\left\|(b(s\,,u(s\,,y))-b(s\,,v(s\,,y))){\bf 1}_{A_{N}(s,y)}\right\|_{k}\mathrm{d}s\,\mathrm{d}y,$
$I_{2}^{2}$ $=8k\int_{0}^{t}\mathrm{d}s\int_{-\infty}^{\infty}\mathrm{d}y\ [p_{t-s}(y-x)]^{2}\left\|(\sigma(s\,,u(s\,,y))-\sigma(s\,,v(s\,,y))){\bf 1}_{A_{N}(s,y)}\right\|_{k}^{2},$
$I_{3}$ $=\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\left\|(b(s\,,u(s\,,y))-b(s\,,v(s\,,y))){\bf 1}_{\Omega\setminus A_{N}(s,y)}\right\|_{k}\mathrm{d}s\,\mathrm{d}y,$
$I_{4}^{2}$ $=8k\int_{0}^{t}\mathrm{d}s\int_{-\infty}^{\infty}\mathrm{d}y\ [p_{t-s}(y-x)]^{2}\left\|(\sigma(s\,,u(s\,,y))-\sigma(s\,,v(s\,,y))){\bf 1}_{\Omega\setminus A_{N}(s,y)}\right\|_{k}^{2},$
where for $N,s>0$ and $y\in\mathbb{R}$,
$A_{N}(s\,,y)=\left\{\omega\in\Omega:\,|u(s\,,y)|(\omega)\leq\mathrm{e}^{N},\ |v(s\,,y)|(\omega)\leq\mathrm{e}^{N}\right\}.$
Observe that the moment bounds obtained in Propositions 4 and 5 only use the linear growth constants $L_{\sigma}$ and $L_{b}$. Therefore, they also hold when $u_{N}$ is replaced by $u$ or $v$. The same is true for the tail estimates obtained Proposition 6, which also hold replacing $u_{N+1}$ by $u$ or $v$. Therefore, we can proceed as in the proof of existence but appeal to the local Lipschitz condition on $b$ and $\sigma$; that is, we recall (2.1) and write
$\left\|\left[b(s\,,u(s\,,y))-b(s\,,v(s\,,y))\right]\mathbf{1}_{A_{N}(s,y)}\right\|_{k}\leq L_{N,b}\left\|u(s\,,y)-v(s\,,y)\right\|_{k},$
and proceed similarly for $\sigma$. Because $\mathrm{P}(\Omega\setminus A_{N}(s\,,y))$ is at most $\mathrm{P}\{|u(t\,,y)|\geq\mathrm{e}^{N}\}+\mathrm{P}\{|v(t\,,y)|\geq\mathrm{e}^{N}\},$ similar computations to those in the proof of existence imply that for all $T>0$,
$\|u(t\,,x)-v(t\,,x)\|_{k}$ $\leq\mathrm{e}^{\beta t}\left[\frac{L_{N,b}}{\beta}+\frac{2\sqrt{k}L_{N,\sigma}}{\beta^{1/4}}\right]\mathcal{N}_{k,\beta,T}(u-v)$
$\quad+\frac{L_{b}C}{256L_{\sigma}^{4}k^{2}}\,\exp\left(512L_{\sigma}^{4}k^{2}t-\frac{N^{3/2}}{2k(512t)^{1/2}L_{\sigma}^{2}}\right)$
$\quad+\sqrt{32k}L_{\sigma}C\mathrm{e}^{512L_{\sigma}^{4}k^{2}t}\exp\left(-\frac{N^{3/2}}{2k(512t)^{1/2}L_{\sigma}^{2}}\right)t^{1/4},$
uniformly for all $\beta>0$, $t\in(0,T)$, $x\in\mathbb{R}$, and $N,k\geq c$, where $C=4(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+1)$ and $c=c(\|u_{0}\|_{L^{\infty}(\mathbb{R})}\,,L_{\sigma}\,,T\,,L_{b})>0$. Then, choosing $\beta=512L_{N,\sigma}^{4}k^{6}$ and $k=\sqrt{L_{N,b}}$, and proceeding as in the proof of existence, we get that
$\adjustlimits{\sup}_{t\in(0,T)}{\sup}_{x\in\mathbb{R}}\|u(t\,,x)-v(t\,,x)\|_{2}\lesssim L_{N,b}^{-1}\exp\left(-\dfrac{N^{3/2}}{2(512T)^{1/2}L_{\sigma}^{2}\sqrt{L_{N,b}}}\right),$
uniformly for all $N\geq c^{\prime}=c^{\prime}(\|u_{0}\|_{L^{\infty}(\mathbb{R})}\,,L_{\sigma}\,,L_{b}\,,T)>0$. The liminf of the right-hand side of the latter expression converges to zero as $N$ tends to infinity [Lemma 7]. Thus, the proof of uniqueness is completed. ∎